Martina leaves home at 9 A.M., bicycling at a rate of 24 mi/h. Two hours later, John leaves, driving at the rate of 48 mi/h. At what time will John catch up with Martina?
Answer by stanbon(57962)

Martina leaves home at 9 A.M., bicycling at a rate of 24 mi/h. Two hours later, John leaves, driving at the rate of 48 mi/h. At what time will John catch up with Martina?
-------------------
Martina DATA:
Rate = 24 mph ; Time = x hrs ; distance = rt = 24x miles
----------------------------
John DATA:
Rate = 48 mph ; Time = x-2 ; distance = rt = 48(x-2) miles
------------------------------
EQUATION:
Martina distance = John distance
24x = 48(x-2)
24x = 48x-96
24x = 96
x = 4 hrs
--------------
John will catch Martina at 9:00 AM = 4 hrs = 1:00 PM
==============
Cheers,
Stan H.